# See More Details: Efficient Image Super-Resolution by Experts Mining

### v1

@@ 4 Experiments
[…]
Datasets and Evaluation.   Following the SR literature (Liang et al., 2021; Chen et al., 2023), we utilize DIV2K (Agustsson & Timofte, 2017) and Flickr2K (Lim et al., 2017) datasets for training. We produce LR images using bicubic downscaling of HR images. When testing our method, we assess its performance on canonical benchmark datasets for SR - Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010), BSD100 (Martin et al., 2001), Urban100 (Huang et al., 2015) and Manga109 (Matsui et al., 2017). We calculate PSNR and SSIM results on the Y-channel from the YCbCr color space.
Implementation Details.   We augment our training data with randomly extracted $64\times 64$-sized crops, with random rotation, horizontal and vertical flipping. Similar to (Sun et al., 2022, 2023), we minimize the L1-Norm between SR output and HR ground truth in the pixel and frequency domain using Adam (Kingma & Ba, 2017) optimizer for $500$K iterations with a batch size of $32$ and initial learning rate of $1\times 10^{-3}$ halving it at following milestones: [$250K$,$400K$,$450K$,$475K$]. All experiments are conducted with the PyTorch framework on NVIDIA RTX 4090 GPUs. We design our smallest model (SeemoRe-T) with 6 RGs. The feature dimension and channel expansion factor in GatedFFN are set to $36$ and $2$, respectively. For all MoRE sub-modules, we select an exponential growth of the channel dimensionality and choose in total of $3$ experts. The kernel size in SEE is set to $11\times 1$ 11. More details can be found in the supplemental, c.f. Table 6.
@@ 4.1 Comparison to State-of-the-Art Methods
Visual Results.   We show visual comparisons (×$\times 4$) in Figure 4. In some challenging scenarios, the previous methods may suffer blurring artifacts, distortions, or inaccurate texture restoration. Contrary to others, our SeemoRe alleviates the blurring artifacts better and maintains structural fidelity. For instance, in image img60 and img73 from Urban100, certain methods like DDistill-SR, SwinIR-Light and DAT-Light fail to accurately reconstruct shadow patterns or window struts, whereas our method exhibits strong recovery of fine details. These visual comparisons highlight SeemoRe’s ability to reconstruct high-quality images by effectively leveraging local and contextual information. Coupled with quantitative comparisons, these findings underscore the effectiveness of our method. More visual results can be found in the Supplementary material.
@@ 4.2 Model Complexity Trade-Off
In the vision domain, scalability becomes more paramount. We strive to expand the limits of our SeemoRe framework, optimizing for both reconstruction fidelity and efficiency. The framework provides three complexity scales—tiny (T), base (B), and large (L)—with progressively improved reconstruction performance, c.f Figure 1. In Table 3, we present comparisons of memory usage and running time, demonstrating that our SeemoRe-T outperforms representative state-of-the-art methods. By using the low-rank feature modulation and simultaneous aggregation of the channel-spatial dependencies, the GPU consumption of our SeemoRe-T is $3\%$ less than DDistill-SR, while being $2$ times faster. Additionally, Table 3 highlights the significant efficiency advantage of SeemoRe over lightweight Transformers. Further results are provided in the Supplemental. To further underscore our method’s capability, we align SwinIR-Light and SRFormer-Light with a size and computational demand similar to ours, followed by retraining these downsized networks using our schedule. The results presented in Table 3 highlight that SeemoRe-T significantly outperforms both Transformer-based models by a considerable margin.
Table 4: Ablation on Blocks. GMACS (normal-$\downarrow$) are computed by upscaling to a $1280\times 720$ HR image.
Method
RME SME Params. GMACS Urban100 Manga109
Baseline
- - 232K 52 31.73 38.63
SeemoRe-T
$\checkmark$ - 249K 48 31.99 38.75
- $\checkmark$ 238K 53 32.05 38.96
@@
[…]
Method
Kernel $k$ Params. GMACS Urban100 Manga109
SeemoRe-T
3 217K 44 32.04 38.84
5 218K 44 32.12 38.72
@@
[…]
Table 5: Ablation on MoRE. Exponential growth yields best performance in terms of parameter counts and PSNR. $\#\mathcal{E}$ denotes the number of experts, and Dim. is the dimensionality of the rank.
@@
$1$ 22{2}2 211K 32.16 38.92
@@ 4.3 Ablation Study
Architecture Contribution.   As reported in Table 4, we evaluate the effectiveness of our proposed key architectural components by comparing them with a baseline model consisting solely of depthwise and pointwise convolutions, more details in Supplemental. After adding the proposed modules into the baseline model, their is a notable and persistent improvement in the results. The incorporation of RME or SME results in improvements of $0.26$ dB or $0.32$ dB on Urban100 over the baseline, respectively. Although both modules individually outperform the baseline with only a marginal increase in parameters, alternating the insertion of both the modules within each RG fully unleashes the model’s capabilities while enhancing the overall efficiency. Overall, our SeemoRe-T obtains a compelling gain of $0.49$ dB and $0.38$ dB on Urban100 and Manga109, respectively.
@@
1234560050505050100100100100Residual Groups##\##Images×$\times 4$ Urban100123456Residual Groups×$\times 4$ Manga109
[…]
@@
Design choices of SME.   The main component of SME module, SEE deploys striped convolutions with large-kernel sizes to effectively module the spatial cues. Table 4 demonstrates that deploying large kernel sizes improves the overall performance of the model. In particular, the PSNR shows a notable gain of 0.18 dB on Urban100 dataset when increasing the kernel size from 3$\times$3 to 11$\times$11 (keeping other settings intact), with only 3K increase in parameters. It clearly proves that such a design benefits in the efficient use of the relevant information to augment the restoration of sharp regions spatially.
Design choices of RME.   We motivate our design choices for the MoRE module in RME by varying the growth function and the number of experts as depicted in Table 5. When pursuing a dynamic solution for determining the optimal low-rank dimensionality, it becomes necessary to design the corresponding search space. First, we present results for ×$\times 2$ upscaling on Urban100 and Manga109 using different growth functions. Based on the observed outcomes, it is evident that an exponentially increasing low-rank dimensionality yields the best performance with marginal increase in the parameters. Hence, we opt to retain this search space design in all further experimentation. Next, we analyze the reconstruction quality based on the number of experts in each MoRE module, while exponentially increasing the low-rank dimensionality. Based on these experiments, we assert that the efficient results are obtained when we have three, as our total number of experts.
@@ 5 Discussion on Experts
Mixture of Low-Rank Experts.   The decision-making process of the router at different network depths is illustrated in Figure 5. Notably, earlier blocks showcase a diverse range of rank choices ( $\mathcal{E}_{1}$,$\mathcal{E}_{2}$,$\mathcal{E}_{3}$), while deeper layers tend to favor lower ranks ($\mathcal{E}_{1}$) (Please note that for every $\mathcal{E}_{i}$, the corresponding rank dimension is $2^{i}$). This phenomenon can be attributed to the hierarchical feature learning nature of deep neural networks, aligning with our expectations. In fact, earlier layers typically capture low-level details and, at times, unwanted noise while reconstructing details in the input LR image, thus resulting in wide variations in the rank choices. In contrast, deeper layers focus on the main structures and key features required for SR. Hence, higher ranks at deeper layers are less favored, as they may introduce redundancy or noise that does not significantly contribute to the overall quality of the reconstructed image. This design aspect provides our method with the flexibility to adapt to the complexity of the task, a capability that, to the best of our knowledge, has not yet been explored in the image reconstruction community. In Figure 5, we further visualize the routing decisions and the corresponding low-rank feature maps for two exemplary input images. It is noteworthy that each individual rank carries distinct information while being mutually complementary. As the model depth increases, the network becomes proficient in restructuring these representations.
How important are MoRE and SEE?   To substantiate the significance of the proposed MoRE and SEE modules, we analyze the feature maps before and after integrating both blocks into the RME module as depicted in Figure 6. This analysis vividly showcases the advantages of leveraging MoRE for contextual information mining within RME. Notably, the activations exhibit reduced noise and enhanced sharpness. Additionally, we observe a synergistic interaction between MoRE and SEE at marked locations (indicated by red arrows): MoRE effectively refines global textures by filtering out noise, while SEE supplements over-filtered regions with critical local details.
@@ 6 Conclusion
We propose a novel ConvNet, named SeemoRe, for efficient and accurate image SR. Our SeemoRe excels in modeling local and contextual information, surpassing both previous CNN-based and lightweight Transformer approaches in terms of efficiency and reconstruction fidelity. Unlike other approaches, we empirically demonstrate both the scalability of efficiency and reconstruction performance. In our approach, we intricately design the rank modulation expert to discern the most pivotal features, enhancing this compressed representation with valuable contextual cues. Our spatial enhancement expert efficiently integrates local spatial-wise information, unlocking the full potential of our architecture. This novel approach optimally exploits the low information regime in the input image, enhancing detail reconstruction while improving efficiency. Extensive experiments on image SR demonstrate that our proposed SeemoRe achieves consistent superior performance over recent state-of-the-art efficient methods on all considered SR benchmarks, while even being on par with the lightweight Transformers in terms of reconstruction fidelity.
7 Impact Statements
@@ Appendix A Further Implementation Details
Here we provide more details about the baseline method for the ablation in Table 4. To ensure roughly equivalent parameter counts and computational complexity, we adopt $5$ RGs with a channel dimensionality of $48$. Within each RG, we integrate simple convolutional operators from our RME submodule without the MoRe module, while the SME module is simplified to a pointwise convolution.
@@ Appendix B More Ablations
Future work and limitations.
The proposed approach, employing a mixture of experts for feature modulation, is versatile for tasks with limited input information, such as low-light enhancement and denoising. Additionally, SeemoRe’s efficient design makes it a valuable solution for dynamic and resource-intensive environments. Expanding the number of experts in our network’s low-rank aspect poses challenges due to rapid feature dimensionality growth. Thus, our approach is currently limited to a small number of experts, contrasting with other fields leveraging larger expert ensembles.
@@
HR Crop
DDistill-SR
ShuffleMixer
SwinIR-Light
DAT-Light
SeemoRe-L
@@
HR Crop
DDistill-SR
ShuffleMixer
SwinIR-Light
DAT-Light
SeemoRe-L
@@
Appendix C Visual Results.

### v2

@@ 4 Experiments
(b) PSNR (dB $\uparrow$) on the Y-Channel. $\ast$ denotes retrained models.
Datasets and Evaluation.
Following the SR literature (Liang et al., 2021; Chen et al., 2023), we utilize DIV2K (Agustsson & Timofte, 2017) and Flickr2K (Lim et al., 2017) datasets for training. We produce LR images using bicubic downscaling of HR images. When testing our method, we assess its performance on canonical benchmark datasets for SR - Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2010), BSD100 (Martin et al., 2001), Urban100 (Huang et al., 2015) and Manga109 (Matsui et al., 2017). We calculate PSNR and SSIM results on the Y-channel from the YCbCr color space.
Implementation Details.
We augment our training data with randomly extracted $64\times 64$-sized crops, with random rotation, horizontal and vertical flipping. Similar to (Sun et al., 2022, 2023), we minimize the L1-Norm between SR output and HR ground truth in the pixel and frequency domain using Adam (Kingma & Ba, 2017) optimizer for $500$K iterations with a batch size of $32$ and initial learning rate of $1\times 10^{-3}$ halving it at following milestones: [$250K$,$400K$,$450K$,$475K$]. All experiments are conducted with the PyTorch framework on NVIDIA RTX 4090 GPUs. We design our smallest model (SeemoRe-T) with 6 RGs. The feature dimension and channel expansion factor in GatedFFN are set to $36$ and $2$, respectively. For all MoRE sub-modules, we select an exponential growth of the channel dimensionality and choose in total of $3$ experts. The kernel size in SEE is set to $11\times 1$ 11. More details can be found in the supplemental, c.f. Table 6.
@@ 4.1 Comparison to State-of-the-Art Methods
Visual Results.
We show visual comparisons (×$\times 4$) in Figure 4. In some challenging scenarios, the previous methods may suffer blurring artifacts, distortions, or inaccurate texture restoration. Contrary to others, our SeemoRe alleviates the blurring artifacts better and maintains structural fidelity. For instance, in image img60 and img73 from Urban100, certain methods like DDistill-SR, SwinIR-Light and DAT-Light fail to accurately reconstruct shadow patterns or window struts, whereas our method exhibits strong recovery of fine details. These visual comparisons highlight SeemoRe’s ability to reconstruct high-quality images by effectively leveraging local and contextual information. Coupled with quantitative comparisons, these findings underscore the effectiveness of our method. More visual results can be found in the Supplementary material.
@@ 4.2 Model Complexity Trade-Off
In the vision domain, scalability becomes more paramount. We strive to expand the limits of our SeemoRe framework, optimizing for both reconstruction fidelity and efficiency. The framework provides three complexity scales—tiny (T), base (B), and large (L)—with progressively improved reconstruction performance, c.f Figure 1. In Table 3, we present comparisons of memory usage and running time, demonstrating that our SeemoRe-T outperforms representative state-of-the-art methods. By using the low-rank feature modulation and simultaneous aggregation of the channel-spatial dependencies, the GPU consumption of our SeemoRe-T is $3\%$ less than DDistill-SR, while being $2$ times faster. Additionally, Table 3 highlights the significant efficiency advantage of SeemoRe over lightweight Transformers. Further results are provided in the Supplemental. To further underscore our method’s capability, we align SwinIR-Light and SRFormer-Light with a size and computational demand similar to ours, followed by retraining these downsized networks using our schedule. The results presented in Table 3(b) highlight that SeemoRe-T significantly outperforms both Transformer-based models by a considerable margin.
Table 4: Ablation on Blocks. GMACS ($\downarrow$) are computed by upscaling to a $1280\times 720$ HR image. We show results for ×$\times 2$ upscaling.
Method RME SME Params. GMACS Urban100 Manga109
Baseline - - 157K 35 31.61 38.55
SeemoRe-T $\checkmark$ - 199K 40 31.96 38.75
- $\checkmark$ 178K 40 31.97 38.87
@@
(a) Contribution of components.
Method Macro Expert Order Urban100 Manga109
SeemoRe-T Spatial - Rank 32.17 38.93
Rank - Spatial 32.22 39.01
(b) Block order.
Method Kernel $k$ Params. GMACS Urban100 Manga109
SeemoRe-T 3 217K 44 32.04 38.84
@@
(c) Kernel size ($k$) variation.
Table 5: Ablation on MoRE. Exponential growth yields best performance in terms of parameter counts and PSNR. $\#\mathcal{E}$ denotes the number of experts and Dim. the rank dimensionality. We show results for ×$\times 2$ upscaling.
@@
$1$ $2$ 211K 32.16 38.92
(a) Low-rank expert design.
Method Recursive Steps $t$ Urban100 Manga109
SeemoRe-T 1 32.10 38.99
2 32.22 39.01
3 31.04 38.13
(b) Recursive step ($t$) variation.
@@ 4.3 Ablation Study
Macro Architecture.
As reported in Table 4(a), we evaluate the effectiveness of our proposed key architectural components by comparing them with a baseline model consisting solely of depthwise and pointwise convolutions, more details in Supplemental. After adding the proposed modules into the baseline model, their is a notable and persistent improvement in the results. The incorporation of RME or SME results in improvements of $0.26$ dB or $0.32$ dB on Urban100 over the baseline, respectively. Although both modules individually outperform the baseline with only a marginal increase in parameters, alternating the insertion of both the modules within each RG fully unleashes the model’s capabilities while enhancing the overall efficiency. Overall, our SeemoRe-T obtains a compelling gain of $0.49$ dB and $0.38$ dB on Urban100 and Manga109, respectively. Moreover, Table 4(b) empirically justifies the chosen block ordering, showcasing the Rank-Spatial macro order design’s superiority over permuted Spatial-Rank macro order. This empirical evidence supplements the qualitative justifications in Section 5 regarding the individual importance of MoRE and SEE blocks.
@@
×$\times 2$ Manga109
1234560050505050100100100100Residual Groups##\##Images×$\times 4$ Urban100
123456Residual Groups×$\times 4$ Manga109
@@
Design choices of SME.
The main component of SME module, SEE deploys striped convolutions with large-kernel sizes to effectively module the spatial cues. Table 4(c) demonstrates that deploying large kernel sizes improves the overall performance of the model. In particular, the PSNR shows a notable gain of 0.18 dB on Urban100 dataset when increasing the kernel size from 3$\times$3 to 11$\times$11 (keeping other settings intact), with only 3K increase in parameters. It clearly proves that such a design benefits in the efficient use of the relevant information to augment the restoration of sharp regions spatially.
Design choices of RME.
We motivate our design choices for the MoRE module in RME by varying the growth function and the number of experts as depicted in Table 5(a). When pursuing a dynamic solution for determining the optimal low-rank dimensionality, it becomes necessary to design the corresponding search space. First, we present results for ×$\times 2$ upscaling on Urban100 and Manga109 using different growth functions. Based on the observed outcomes, it is evident that an exponentially increasing low-rank dimensionality yields the best performance with marginal increase in the parameters. Hence, we opt to retain this search space design in all further experimentation. Next, we analyze the reconstruction quality based on the number of experts in each MoRE module, while exponentially increasing the low-rank dimensionality. Based on these experiments, we assert that the efficient results are obtained when we have three, as our total number of experts. Further, we ablate the choice of recursive steps for SeemoRe-T in Table 5(b), where our plain version takes ($t=2$). It can be seen that lower ($t=1$) and higher ($t=3$) values either fail to capture sufficient contextual information or overly compromise spatial image features.
@@ 5 Discussion on Experts
Mixture of Low-Rank Experts.
The decision-making process of the router at different network depths is illustrated in Figure 5(a). Notably, earlier blocks showcase a diverse range of rank choices ( $\mathcal{E}_{1}$,$\mathcal{E}_{2}$,$\mathcal{E}_{3}$), while deeper layers tend to favor lower ranks ($\mathcal{E}_{1}$) (Please note that for every $\mathcal{E}_{i}$, the corresponding rank dimension is $2^{i}$). This phenomenon can be attributed to the hierarchical feature learning nature of deep neural networks, aligning with our expectations. In fact, earlier layers typically capture low-level details and, at times, unwanted noise while reconstructing details in the input LR image, thus resulting in wide variations in the rank choices. In contrast, deeper layers focus on the main structures and key features required for SR. Hence, higher ranks at deeper layers are less favored, as they may introduce redundancy or noise that does not significantly contribute to the overall quality of the reconstructed image. This design aspect provides our method with the flexibility to adapt to the complexity of the task, a capability that, to the best of our knowledge, has not yet been explored in the image reconstruction community. In Figure 5(b), we further visualize the routing decisions and the corresponding low-rank feature maps for two exemplary input images. It is noteworthy that each individual rank carries distinct information while being mutually complementary. As the model depth increases, the network becomes proficient in restructuring these representations.
How important are MoRE and SEE?
To substantiate the significance of the proposed MoRE and SEE modules, we analyze the feature maps before and after integrating both blocks into the RME module as depicted in Figure 6. This analysis vividly showcases the advantages of leveraging MoRE for contextual information mining within RME. Notably, the activations exhibit reduced noise and enhanced sharpness. Additionally, we observe a synergistic interaction between MoRE and SEE at marked locations (indicated by red arrows): MoRE effectively refines global textures by filtering out noise, while SEE supplements over-filtered regions with critical local details.
@@ 6 Conclusion
We propose a novel ConvNet, named SeemoRe, for efficient and accurate image super-resolution. Our SeemoRe excels in modeling local and contextual information, surpassing both previous CNN-based and lightweight Transformer approaches in terms of efficiency and reconstruction fidelity. Unlike other approaches, we empirically demonstrate both the scalability of efficiency and reconstruction performance. In our approach, we intricately design the rank modulation expert to discern the most pivotal features, enhancing this compressed representation with valuable contextual cues. Our spatial enhancement expert efficiently integrates local spatial-wise information, unlocking the full potential of our architecture. This novel approach optimally exploits the low information regime in the input image, enhancing detail reconstruction while improving efficiency. Extensive experiments on image super-resolution demonstrate that our proposed SeemoRe achieves consistent superior performance over recent state-of-the-art efficient methods on all considered SR benchmarks, while even being on par with the lightweight Transformers in terms of reconstruction fidelity.
Impact Statement
@@
Acknowledgments
This work was supported by The Alexander von Humboldt Foundation.
@@ Appendix A Further Implementation Details
Here we provide more details about the baseline method for the ablation in Tables 4(a) and 7. In the main text, we evaluate SeemoRe-T by sequentially removing the proposed RME and SME blocks, resulting in a plain baseline model with fewer parameters and GMACs. To ensure a fair comparison, we adjust the baseline configuration to match our plain SeemoRe-T model. To ensure roughly equivalent parameter counts and computational complexity, we adopt $5$ RGs with a channel dimensionality of $48$. Within each RG, we integrate simple convolutional operators from our RME submodule without the MoRE module, while the SME module is simplified to a pointwise convolution.
Comparison to lightweight SR Transformers (×$\times 3$).
In Table 8, we present the performance of our SeemoRe-L model for ×$\times 3$ upscaling, extending the results from Table 2 in the main text. Our SeemoRe-L consistently outperforms other lightweight Transformers and demonstrates only a slightly lower performance compared to DAT-Light (Chen et al., 2023).
@@ Appendix B More Ablations
B.1 Architecture Design
SEE block compared to prior designs.
The results in Table 10 prove that our SEE block design outperforms the FusedMB-Conv block proposed by ShuffleMixer (Sun et al., 2022) in terms of reconstruction abilities while maintaining higher efficiency. Moreover, substituting the large-kernel convolution in the Conv2Former (Hou et al., 2022) block with our striped large-kernel variant not only enhances efficiency but also improves the reconstruction capabilities of high-frequency information, as evident from Urban100 results.
MoRE block design.
Our rationale behind the MoRE design involves the aggregation of valuable contextual information. Similar to prior works (Liu et al., 2020b), we assign more learning parameters to enhance the high-frequency features while keeping the simple DConv-branch as residual to facilitate the optimization. We further support this rationale with empirical evidence provided in Table 11. The results show that adding the extended feature to the output of DConv performs better than with and without the aggregation output.
@@ B.1 Architecture Design
Futhermore, we investigate increasing the number of experts to $8$ as shown in Table 14 the impact on the overall model performance. The results indicate that increasing the number of experts adds complexity, however it doesn’t consistently improve the reconstruction fidelity. Balancing the low-rank space and the expert count offers to fine-tune the performance trade-off. Though, our emphasis here is on efficiency, we aim to explore more complex designs in future research.
B.2 Evaluation on Real SR
We conduct experiments for Real SR (×$\times 4$) using the Real-ESRGAN (Wang et al., 2021) degradation model on SeemoRe-T and the current efficient SOTA SR model SAFMN (Sun et al., 2023), see Table 15. Both SAMFN and SeemoRe-T are initialized from the ×$\times 4$ bicubic checkpoints, we reduce the number of iterations on the DF2K_OST dataset by half (250k) and train only using the L1 loss. We report the popular NR-IQA metrics (NIQE and BRISQUE) on the commonly used real-world image collection given in SwinIR (Liang et al., 2021). Additionally, we conduct a cross-dataset evaluation using testsets with more realistic degradation of different severity levels (Type I and Type II), as provided by (Liang et al., 2022).
Table 15: Real SR performance. NIQE and BRISQUE are reported on the real image collection provided by SwinIR (Liang et al., 2021). DIV2K-I and DIV2K-II performance reported as PSNR.
Method NIQE ($\downarrow$) BRISQUE ($\downarrow$) DIV2K-I DIV2K-II
Bicubic 7.65 58.29 26.30 25.71
SAFMN 7.19 51.39 26.80 26.77
SeemoRe-T 6.53 45.53 27.07 27.01
Appendix C Future work and limitations
The proposed approach, employing a mixture of experts for feature modulation, is versatile for tasks with limited input information, such as low-light enhancement and denoising. Additionally, SeemoRe’s efficient design makes it a valuable solution for dynamic and resource-intensive environments. Expanding the number of experts in our network’s low-rank aspect poses challenges due to rapid feature dimensionality growth. Thus, our approach is currently limited to a small number of experts, contrasting with other fields leveraging larger expert ensembles. Despite the improving trade-off between efficiency and reconstruction fidelity, as depicted in Figures 8 and 7, our SeemoRe model still contends with blur artifacts. However, similar artifacts can also be observed in Transformer-based super-resolution alternatives, albeit at a higher computational cost (in terms of inference time). While our model represents a pioneering effort in utilizing a mixture of low-rank experts for super-resolution, significant opportunities for further research exist. For instance, exploring explicit constraints on the features learned by different experts presents intriguing research directions with potential applications across a spectrum of restoration problems. We wish our network serve as a straightforward yet effective baseline, stimulating continued exploration in the field.
@@ Appendix C Future work and limitations
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
HR Crop DDistill-SR ShuffleMixer SwinIR-Light DAT-Light SeemoRe-L
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
HR Crop DDistill-SR ShuffleMixer SwinIR-Light DAT-Light SeemoRe-L
@@
Appendix D Visual Results.